OpenStudy (anonymous):

Let T1 : R3 −→ R2 be the linear transformation given by the matrix [ 1 3 2] [2 0 −1]  and let T2 : R2 −→ R1 be the linear transformation given by the matrix  [5 3] The kernel of T1 is one dimensional. Find a vector u1 which spans Ker(T1) (i.e., a basis for Ker(T1)). 

OpenStudy (blacksteel):

Let vector v = defined as \[a\\b\\c\] Then T1(v) =\[a+3b+2c\\2a-c\] The kernel of a transformation T is defined as the set of all vectors k for which T(k)=0 Hence, the kernel of T1 will be all vectors v such that a + 3b + 2c = 0 2a - c = 0 Let's use a as the free variable. From the second equation, we get c = 2a. From the first equation we get a + 3b + 2c = 0 a + 3b + 4a = 0 5a + 3b = 0 3b = -5a b = (-5/3)a Then the kernel of T1 is all vectors v of the form \[a\\(-5/3)a\\2a\] Since the basis is one-dimensional, any vector satisfying this form will span Ker(T1) and thus be a valid choice for u1, so pick one. For example, let a = 1. Then u1=\[1\\-5/3\\2\]

OpenStudy (blacksteel):

(You can plug the vector u1 into the transformation and see if you get the 0 vector to test whether it's in the kernel)

OpenStudy (anonymous):

wow thank you so so so much